Today Janet's age in years is three times her dog's age ing months, In one year, the dog's age in months will be 2 times Janet's age then, in years. Today Janet is _________ years old. Please show your work.
Answer by ankor@dixie-net.com(22740)

Today Janet's age in years is three times her dog's age in months,
In one year, the dog's age in months will be 2 times Janet's age then, in years.
:
Let j = Janet's age in years
Let d = dog's age in months
:
"Janet's age in years is three times her dog's age in months,"
j = 3d
:
"In one year, the dog's age in months will be 2 times Janet's age then, in years."
d + 12 = 2(j+1)
d + 12 = 2j + 2
d = 2j + 2 - 12
d = 2j - 10
replace j with 3d
d = 2(3d) - 10
10 = 6d - d
10 = 5d
d = 10/5
d = 2 months is the dogs age
then
j = 3(2)
j = 6 yrs is Janet's age
:
:
Check this in the statement:
"In one year, the dog's age in months will be 2 times Janet's age then, in years.
2 + 12 = 2(6+1)
14 = 2(7)